How much ounces of vegetable oil containing 8% saturated fats must be mixed with 25 ounces of solution containing 90% olive oil to blend a salad dressing which is 54% olive oil?
Answer by jorel1380(3719)

25x.9=22.5
.54p=22.5
p=22.5/.54=41-2/3
(41-2/3)-25=16-2/3
You need 16-2/3 ozs. of vegetable oil for your salad dressing..